OpenStudy (anonymous):

What is 3/2* 3/5* 7/8 equal to in fraction form.

OpenStudy (lgbasallote):

hint: \[\huge \frac 32 \times \frac 35 \times \frac 78 \implies \frac{3 \times 3 \times 7}{2 \times 5 \times 8}\] does that help?

OpenStudy (mayankdevnani):